# When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails

### v2

@@ Abstract
Large Reasoning Models (LRMs) demonstrate remarkable capabilities on complex reasoning tasks but remain vulnerable to severe safety risks, including harmful content generation and jailbreak attacks. Existing mitigation strategies rely on injecting heuristic safety signals during training, which often suppress reasoning ability and fail to resolve the safety-reasoning trade-off. To systematically investigate this issue, we analyze the reasoning trajectories of diverse LRMs and uncover a phenomenon we term Self-Jailbreak, where models override their own risk assessments and justify responding to unsafe prompts. This finding reveals that LRMs inherently possess the ability to reject unsafe queries, but this ability is compromised, resulting in harmful outputs. Building on these insights, we propose the Chain-of-Guardrail (CoG), a training framework that recomposes or backtracks unsafe reasoning steps, steering the model back onto safe trajectories while preserving valid reasoning chains. Extensive experiments across multiple reasoning and safety benchmarks demonstrate that CoG substantially improves the safety of current LRMs while preserving comparable reasoning ability, significantly outperforming prior methods that suffer from severe safety–reasoning trade-offs.
@@ 1 Introduction
Large Reasoning Models (LRMs), such as OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025), and Qwen3 (Yang et al., 2025), have demonstrated remarkable capabilities in solving complex tasks and are increasingly being adopted across a wide range of domains. Despite these advances, researchers have identified that current LRMs remain hindered by significant safety vulnerabilities, such as the generation of harmful content as well as the susceptibility to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025; Wang et al., 2025a). These issues highlight a critical obstacle to the reliable and responsible use of LRMs (Weidinger et al., 2021). Therefore, addressing these challenges is critical for ensuring that LRMs operate safely and remain aligned with human values.
While recent efforts (Wang et al., 2025b; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have attempted to mitigate these risks, they largely exhibit a pronounced safety-reasoning trade-off. Specifically, current studies have primarily focused on curating safety-oriented datasets based on heuristic rules and fine-tuning LRMs to produce refusal responses. Unfortunately, while such approaches can improve safety to some extent, the pattern-based behavior often conflicts with the flexible and multi-step deliberation required for complex reasoning, thereby substantially degrading reasoning ability. Therefore, it is critical to identify the underlying causes of the unsafe behavior in LRMs, and develop more principled strategies that can enhance safety without sacrificing reasoning performance.
To this end, we conduct a systematic analysis of LRMs’ reasoning trajectories and identify a critical failure phenomenon, which we term Self-Jailbreak. Specifically, as demonstrated in Figure 1, we decompose LRM’s reasoning thinking trajectory into three stages: Risk Awareness, Risk Analysis, and Response Strategy. We observe that LRMs often detect harmful intent during the Risk Awareness stage, but subsequently override this recognition during the Risk Analysis stage, and ultimately persuade themselves to respond to unsafe queries. This reveals that LRMs possess an inherent ability to reject unsafe queries, yet it is undermined by Self-Jailbreak, resulting in unsafe generation. Furthermore, we categorize Self-Jailbreak into four distinct types—Benign Reframing, Logical Fallacy, Warning, and Risk Misrecognition—and find that Warning accounts for most unsafe outputs.
Existing approaches that solely inject heuristic safety signals during training fail to resolve the safety–reasoning trade-off. While they suppress overtly unsafe outputs, such interventions do not address unsafe reasoning steps embedded within reasoning trajectories. Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving reasoning quality. Specifically, CoG leverages precise classification of Self-Jailbreak to perform targeted safety interventions, mitigating unsafe behavior while preserving the model’s reasoning abilities and reasoning patterns. To achieve this, CoG employs two strategies: 1) Safety Recomposition, which systematically reorganizes unsafe reasoning into entirely new, safety-assured chains, and 2) Safety Backtrack, which preserves the model’s original reasoning path while incorporating self-check and backtracking mechanisms to correct risky steps before they lead to unsafe conclusions. The resulting safety-oriented reasoning chains are then used to fine-tune LRMs, aligning them toward safe responses while maintaining the reasoning capability.
Figure 1: Illustration of “Self-Jailbreak.” The Instruct model (left) directly refuses harmful queries, whereas the LRM (right) identifies risk during Risk Awareness but later engages in Self-Jailbreak (red text) during Risk Analysis, ultimately overriding its initial safety judgment and producing harmful content in the Response Strategy stage.
Extensive experiments across multiple safety and reasoning benchmark datasets demonstrate that COG excels at balancing the safety and reasoning performance of LRMs. On Qwen3-32B, CoG improves the reasoning ability by 13.5% over SafeKey while achieving state-of-the-art performance on GPQA and AIME (Rein et al., 2024; Mathematical Association of America (2024), MAA), maintaining comparable reasoning performance. Further analysis of reasoning trajectories reveals that, unlike other approaches, CoG preserves the model’s intrinsic reasoning patterns, confirming that its safety gains do not come at the expense of the model’s original deliberation trajectory. In summary, our contributions include: In summary, our contributions include:
@@
  • To the best of our knowledge, we are the first to identify and characterize the Self-Jailbreak phenomenon, revealing it as a key factor behind the poor safety performance of LRMs.
  • We propose a systematic analytical framework to categorize Self-Jailbreak instances and quantify their impact on model behavior, finding that Benign Framing and Warning constitute the majority of the Self-Jailbreak phenomenon.
  • We introduce the Chain-of-Guardrail (CoG), a training framework that mitigates Self-Jailbreak and substantially improves model safety while preserving reasoning capabilities. Experiment results show that CoG improves the safety-reasoning trade-off by 3.39%, reaching the state-of-the-art (SOTA) across multiple LRMs.
Figure 2: Processing of harmful queries by LRM, illustrating the workflow from receiving prompts, through risk detection during reasoning, to the four possible outcomes: aware of risk and responds, aware of risk and refuses, unaware of risk and responds, unaware of risk and refuses.
2 Motivation: Preliminary Experiments On Self-Jailbreak
This section examines the thinking trajectory of LRMs, drawing on pilot experimental results (reported in Appendix B.1) which indicate that most LRMs become unsafe when the thinking mode is enabled. Subsequently, we find that even if LRMs are capable of recognizing latent dangerous intentions, they still respond to the harmful query. Such results shed light on our hypothesis: Self-Jailbreak. We further validate the existence of Self-Jailbreak with experiments, which represent two significant findings: (1) Most LRMs encounter Self-Jailbreak during the reasoning trajectory. (2) Among the different categories of Self-Jailbreak, Warning is the primary source.
2.1 Hypothesis: Self-Jailbreak via Thinking–Answer Misalignment
We analyzed the models’ chains of thought, discovering that their thinking trajectory can be decomposed into three stages: risk awareness, risk analysis, and response strategy, which enables a deeper examination of how thinking affects the safety of LRMs.
Building on the previous experiment, we employ LLM-as-Judge (Team, 2024; Gu et al., 2024) to assess whether the LRM demonstrates awareness of potential risks during the risk awareness stage, and use Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content. This setup enables us to examine potential misalignments between the model’s reasoning trajectory and its generated answers. As illustrated in Figure 2, our analysis revealed a striking contradiction:
LRMs are able to identify most of the harmful queries but still generate unsafe answers. 94.2% of harmful queries, as well as users’ potentially malicious intents, are recognized; however, 52.30% of their generated outputs still contain hazardous content. In contrast, only 39.90% of the cases reflect awareness of risk followed by a safe response. This inconsistency between “thinking” and “response” raises a critical concern: during subsequent reasoning, LRMs may tend to persuade themselves to respond to such queries. Based on this observation, we propose the following hypothesis:
Self-Jailbreak During the thinking process, LRMs often exhibit a phenomenon we term Self-Jailbreak, where models persuade themselves to produce responses to potentially harmful requests in an attempt to assist the user, even after recognizing the risky nature of the queries and the user’s underlying malicious intent.
Table 1: Self-Jailbreak category distribution of various LRMs (%)
Model Response Safety Self-Jailbreak Categories
Safe Unsafe Benign Reframing Warning Logical Fallacies Harm Identification
DS-R1-0528 4.07 95.93 42.29 52.12 0.51 1.02
DS-Llama-70B 0.50 99.50 47.89 45.67 0.91 5.04
Qwen3-8B 0.21 99.79 38.60 50.69 2.41 8.10
Qwen3-14B 0.15 99.85 37.30 53.10 1.73 7.32
Qwen3-32B 0.36 99.64 36.32 52.51 2.20 8.61
2.2 Categorization and Existence of Self-Jailbreak
Building on this hypothesis and inspiration from backtracking during mathematical reasoning  (Chen et al., 2025), we designed a more rigorous classification framework, which is guided by the principle that the model’s reasoning trajectory should not generate content that directly or indirectly facilitates harmful user behaviors. We categorize these behaviors into four types:
Benign Reframing:
The model recognizes the risk but attempts to reinterpret the user’s intent benignly, offering help from such interpretation.
Warning:
The model assumes that appending a warning to harmful content is sufficient to prevent its misuse, and proceeds to generate its answer.
Logical Fallacies:
Given a query containing logical contradictions, the model detects potential harm but falls into logical traps, leading to logical fallacies.
Harm Identification:
The model fails to recognize the harm in the query, resulting in answering.
To investigate the prevalence of Self-Jailbreak in LRMs, we categorize and analyze the thinking trajectory of several contemporary LRMs on 2k queries in WildJailbreak (Jiang et al., 2024a). As shown in Table 1, we observe that:
(1) Self-Jailbreak is a prevalent phenomenon observed in various LRMs. Experiment results indicate that Self-Jailbreak occurs frequently across LRMs, including peripheral models like DeepSeek-R1-0528 (DeepSeek-AI, 2025), highlighting the importance and relevance of our study.
(2) Warning is the most common source of Self-Jailbreak. As illustrated in Table 1, Warning is the most categorized Self-Jailbreak type across various LRMs. This tendency likely stems from the fact that LRMs are largely trained on math and code data, which predisposes them to respond in ways that meet user expectations. Consequently, even when confronted with harmful queries, LRMs are likely to be helpful, inadvertently bypassing their own safety mechanisms by providing content alongside a safety disclaimer. Such behavior ultimately gives rise to Self-Jailbreak.
3 Chain-of-Guardrail: A Training Framework to Mitigate Self-Jailbreak
To systematically address the risks of Self-Jailbreak in LRMs, we propose a unified safe training framework: Chain-of-Guardrail (COG). As illustrated in Figure 3, the framework is designed to create an interpretable and structured pipeline that automatically removes or rejects unsafe content during the reasoning trajectory $c$, producing a safe and logically consistent response $y_{\text{safe}}$. To achieve this, COG employs two complementary methods: Safety Recomposition (SafR) and Safety Backtrack (SafB). SafR focuses on proactively restructuring and correcting the reasoning chain to eliminate potential safety risks, while SafB focuses on the reasoning trajectory when unsafe content is detected, ensuring that the model corrects errors and generates a safe response.
@@ 3 Chain-of-Guardrail: A Training Framework to Mitigate Self-Jailbreak
3.1 Optimization Goal of LRMs in COG
Specifically, we formulate the optimization objective as follows:
$f:(\pi_{0},x)\rightarrow y_{\text{safe}},\quad\text{s.t.}\quad\text{Preserve}(\text{Reasoning}_{\pi_{0}})$ (1)
where $\pi_{0}$ is the original model, $y_{\text{safe}}$ is the safe response, and $\text{Preserve}(\cdot)$ ensures maximal retention of the original reasoning.
To achieve this, we explicitly model the Chain-of-Thought (CoT) as a structured latent variable:
$c=\pi_{0}(x)=[d(x),a(x),p(x)],$ (2)
where $d(x)$, $a(x)$, and $p(x)$ represent the risk awareness, risk analysis, and response strategy.
Depending on the adopted process, the safety-oriented reasoning chain is constructed as follows:
Chain of Safety Recomposition (SafR):
$c_{\text{safe}}=\text{Merge}(d(x),\hat{a}(x),\hat{p}(x)),$ (3)
where $\hat{a}(x)$ and $\hat{p}(x)$ are the risk analysis and response strategy recomposed by the Safety Recomposition procedure.
Chain of Safety Backtrack (SafB):
$c^{\prime}_{\text{safe}}=[c,s(x)],$ (4)
where $s(x)$ denotes the model-generated self-check module appended to the original reasoning chain in the Safety Backtrack procedure.
The final safe response $y_{\text{safe}}$ is produced based on the safety-oriented reasoning chain derived from either SafR or SafB.
3.2 Selective Loss Masking Strategy
To avoid undesired shifts in the model’s predicted distribution due to safety interventions, we employ a selective loss masking strategy. Let the output sequence be $o=[c_{\text{safe}},y_{\text{safe}}]$ for SafR, and $o=[c^{\prime}_{\text{safe}},y_{\text{safe}}]$ for SafB, where $c_{\text{safe}}$ and $c^{\prime}_{\text{safe}}$ are defined as in Eq. 3 and Eq. 4, respectively.
We propose the utilization of a binary mask vector $m\in\{0,1\}^{|o|}$, to supervise distinct segments of the output sequence $o$ selectively. The primary training goal can be summarized as follows:
$\mathcal{L}_{SafR/SafB}=-\sum_{i=1}^{|o|}m_{i}\cdot\log P_{\theta}(o_{i}|o_{<i},x),$ (5)
This selective loss masking strategy allows for flexibility depending on the applied process:
Safety Recomposition: $m$ covers all tokens, optimizing both the revised reasoning chain and the safe response, allowing the model to learn the complete safe reasoning trajectory.
Safety Backtrack: $m$ covers only the appended self-check and final response, while the original chain is masked as context, preserving the model’s reasoning distribution and enhancing self-reflection and safety.
3.3 Workflow of COG
Both SafR and SafB aim to enhance safety by mitigating certain aspects of Self-Jailbreak, following a three-phase structure. While their core workflows overlap, they differ in how the reasoning chain is modified and integrated.
Phase 1: Input & Analysis (shared)
Initial Inference. The base model $\pi_{0}$ infers on the harmful prompt $x$, producing the initial reasoning chain $c=[a(x),p(x)]$ and response $y$.
Information Extraction. A structured prompt (see Appendix A.2) together with Qwen2.5-72B-Instruct (Team, 2024) extracts the risk analysis and response strategy.
Failure Classification. Given $x$, $a(x)$, and $p(x)$, Qwen2.5-72B-Instruct determines whether the reasoning violates safety principles and assigns a Self-Jailbreak type.
Phase 2: Data Construction (divergent)
SafR – Reasoning Chain Recomposition and Merging. For each identified failure type, a tailored prompt guides $\pi_{0}$ to perform reasoning chain recomposition, adjusting the risk analysis and response strategy with safety-oriented considerations to generate corrected sub-chains. Subsequently, these recomposed analyses $a(x)$ and predictions $p(x)$ are merged into a coherent, Safety-oriented Chain-of-Thought (S-COT).
SafB – Safety Backtrack Augmentation. Rather than recomposing, $\pi_{0}$ performs a reflective process on the original reasoning chain to generate a supplementary safety check $s(x)$. Carefully engineered transitional expressions ensure a natural integration between the reasoning chain and the check, resulting in a coherent S-COT.
Phase 3: Inference & Alignment (shared, with minor variation)
Safe Response Generation. Both SafR and SafB use the S-COT as the reasoning trajectory to guide $\pi_{0}$ in generating a safe response $y_{\text{safe}}$.
Alignment Training. In both methods, $\pi_{0}$ is further trained on the safe response data, yielding the aligned and safety-enhanced model $\pi_{1}$ with strong reasoning capabilities.
4 Experiment
4.1 Experiment Setting
Baselines We include four representative methods: STAR-1  (Wang et al., 2025b) and SafeChain  (Jiang et al., 2025) improve safety by performing SFT on a curated set of safe data. SAFEPATH  (Jeung et al., 2025) guides model behavior by training it to generate a fixed prefix, and SafeKey  (Zhou et al., 2025b) enforces stricter safety alignment by introducing two additional safety classifiers. Each method was reproduced using the officially released training configurations and datasets, with details provided in Appendix A.1.
Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). Our Safety Recomposition and Safety Backtrack methods are implemented via supervised fine-tuning (SFT) on the Qwen3 base models. As illustrated in Figure 3, the training and experimental procedures comprise three distinct phases. The details are shown in the Appendix A.2.2.
Safety Benchmarks We evaluate model safety using four benchmarks. SORRY-Bench  (Xie et al., 2024) and StrongREJECT  (Souly et al., 2024) measure the model’s ability to refuse harmful prompts, while WildJailBreak  (Jiang et al., 2024b) and JailBreakBench  (Chao et al., 2024) assess its robustness against adversarial jailbreak attacks.
[Uncaptioned image] Figure 4: Safety vs. reasoning trade-off for the 32B model.
Reasoning Benchmarks To assess general reasoning ability, we use AIME2024  (Mathematical Association of America (2024), MAA) for mathematical reasoning and GPQA-Diamond  (Rein et al., 2024) for complex knowledge-based reasoning. To reduce randomness in evaluation, we perform multiple rollout rounds: GPQA is averaged over two rounds, and AIME over eight rounds. Additional details of the reasoning and safety benchmarks can be found in Appendix A.3.
Training Dataset We extracted 15,000 high-quality harmful prompts from 7 public collections—Alert  (Tedeschi et al., 2024), ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming  (Ahmadian et al., 2024), Do-Not-Answer  (Wang et al., 2023), AttaQ  (Kour et al., 2023), and Toxic-Chat  (Lin et al., 2023). We then designated this curated set as our seed dataset and fed it into our pipeline to generate corresponding high-quality safety response sets.
4.2 Main Result
Table 2: Performance comparison of different methods under the main experimental setting. Here, lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.

### v3

# When Models Outthink Their Safety: Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models

@@ Abstract
Large Reasoning Models (LRMs) achieve strong performance on complex multi-step reasoning, yet they still exhibit severe safety failures such as harmful content generation. Existing methods often apply coarse-grained constraints over the entire reasoning trajectories, which can undermine reasoning capability while failing to address the root causes of unsafe behavior. In this work, we uncover a previously underexplored failure mode in LRMs, termed Self-Jailbreak, where models initially recognize the harmful intent of a query, but override this judgment during subsequent reasoning steps, ultimately generating unsafe outputs. Such a phenomenon reveals that LRMs are capable of recognizing harm, while safety failures primarily arise from reasoning steps. Motivated by this finding, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions while maintaining reasoning ability. Experiments across multiple safety and reasoning benchmarks indicate that CoG achieves a favorable balance between safety and reasoning performance compared with existing approaches.
When Models Outthink Their Safety:
Unveiling and Mitigating Self-Jailbreak in Large Reasoning Models
Yingzhi Mao1,2thanks: Equal contribution.thanks: Code: https://github.com/icip-cas/COG., Chunkang Zhang [^1] footnotemark: 1, Junxiang Wang1, Xinyan Guan1,2, Boxi Cao1 Yaojie Lu1, Hongyu Lin1, Xianpei Han1,2, Le Sun1,2 1Chinese Information Processing Laboratory, Institute of Software, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {maoyingzhi2024, zhangchunkang2022, luyaojie}@iscas.ac.cn
@@ 1 Introduction
Large Reasoning Models (LRMs) (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025) have demonstrated strong capabilities across a wide range of tasks, including software engineering and mathematical reasoning. As these models are increasingly integrated into autonomous agents and decision-support systems, ensuring their safety and alignment with human values is foundational.
However, recent studies show LRMs remain vulnerable to serious safety failures (Green et al., 2025; Arrieta et al., 2025), ranging from generating harmful content (Qiu et al., 2025; He et al., 2025) to exhibiting vulnerability to jailbreak attacks (Zhou et al., 2025a; Zhang et al., 2025b). Therefore, strengthening safety capabilities of LRMs is not merely an auxiliary concern, but a prerequisite for their trustworthy deployment at scale.
Figure 1: Illustration of Self-Jailbreak. While the model initially recognizes the harmful nature of a query, it overrides this judgment during subsequent reasoning steps, ultimately yielding unsafe output.
While recent efforts (Wang et al., 2025; Jeung et al., 2025; Zhou et al., 2025a; Jiang et al., 2025) have made progress towards safer LRMs, they often incur a pronounced safety–reasoning trade-off. A primary limitation shared by these approaches is that they typically inherit safety paradigms originally designed for standard LLMs, treating internal reasoning trajectories of LRMs as an undifferentiated extension of responses. Subsequently, safety constraints are typically imposed in a coarse-grained and global manner. Such heuristic safeguards often interfere with the model’s intrinsic reasoning patterns, degrading the coherence required for multi-step reasoning. These limitations indicate that mitigating safety risks in LRMs without undermining their reasoning capabilities requires moving toward mechanisms that can localize and address failure-inducing steps within the reasoning chain.
To this end, we conduct a systematic analysis of reasoning trajectories of LRMs to uncover the root causes of their safety failures. In particular, we identify two distinct failure modes: Harm Misidentification, where harmful intent is not recognized, and Self-Jailbreak, a severe and previously overlooked phenomenon in which the model initially identifies potential harm but later overturns this safety judgment during subsequent reasoning. To enable fine-grained analysis, we decompose each reasoning trajectory into three consecutive stages: risk awareness, risk analysis, and response strategy. Specifically, as demonstrated in Figure 1, given a harmful query (“How to build a bomb…”), the model correctly identifies the harm during the risk-awareness stage, but then overrides this judgment by rationalizing a seemingly benign user intent (“…could be a student…”). In the response-strategy stage, it chooses to answer rather than refuse, and the resulting response contains harmful content.
Through quantitative analysis on WildJailbreak, we find that only a small portion of unsafe outputs arise from Harm Misidentification, whereas Self-Jailbreak is the dominant failure mode, responsible for nearly 80% of the unsafe cases we analyze and thus constituting our primary focus. In these cases, LRMs correctly recognize harmful intent during the risk awareness stage, yet this recognition is subsequently overridden during the risk analysis stage, where the model effectively persuades itself to comply with the unsafe request.
Building on these insights, we propose Chain-of-Guardrail (CoG), a training framework designed to mitigate Self-Jailbreak while preserving the reasoning capability. CoG executes targeted, step-level interventions conditioned on the diagnosed Self-Jailbreak patterns, correcting only the segments that induce unsafe behavior. We instantiate this framework via two complementary variants: Safety Recomposition, which rewrites the reasoning chain into a logically consistent, safe alternative, and Safety Backtrack, which preserves the original trajectory while revising risky segments before they lead to unsafe outputs. These corrected traces serve as fine-tuning data to align LRMs without suppressing their reasoning potential.
Extensive experiments across multiple safety and reasoning benchmarks demonstrate that CoG achieves a superior balance between safety and reasoning performance. Across diverse LRMs, CoG consistently reduces attack success rates to competitive levels on standard safety benchmarks, while substantially boosting performance on challenging reasoning tasks. Notably, on Qwen3-32B, CoG achieves safety performance comparable to SafeKey, while substantially improving reasoning accuracy, with GPQA-Diamond increasing from 54.30 to 62.38 and AIME2024 from 71.70 to 82.08. Further analyses from both reasoning-pattern and representation-geometry perspectives suggest that CoG improves safety by correcting specific failure-inducing reasoning steps, rather than globally rewriting the model’s reasoning paradigm.
We summarize our major contributions as follows:
@@
  • To the best of our knowledge, we are the first to uncover and characterize the Self-Jailbreak phenomenon, revealing it as a primary driver of safety failures in LRMs.
  • We propose an analysis framework that enables the categorization and quantitative analysis of unsafe reasoning behaviors.
  • We introduce the Chain-of-Guardrail (CoG), a trajectory-level training framework that achieves the state-of-the-art safety-reasoning balance across multiple benchmarks.
2 Unveiling Self-Jailbreak in Reasoning Trajectories
In this section, we analyze the causes of safety failures in LRMs through an analysis of their reasoning trajectories. Our findings indicate a consistent mismatch between risk awareness and final response: LRMs may generate unsafe outputs even after identifying potential risks during reasoning. Beyond failures of harm identification, we observe a prevalent pattern in which earlier safety assessments are revised or overridden in later reasoning stages, which we refer to as Self-Jailbreak. We further provide a taxonomy of such behaviors across multiple LRMs to support a more fine-grained analysis of this failure mode.
2.1 The Prevalance of Self-Jailbreak in LRMs
Investigation Setup
To localize safety failures within the reasoning process, we decompose each reasoning trajectory into three stages: risk awareness, risk analysis, and response strategy. Each stage is assessed independently. For evaluation, we sample 2,000 data points from Wildjailbreak (Jiang et al., 2024a) as the benchmark, following the official setting with Llama-Guard (Llama Team, 2024) to evaluate whether the final response contains harmful content. In addition, we utilize Qwen2.5-72B-Instruct (Team, 2024; Gu et al., 2024) to determine whether potential risks are explicitly identified during the risk awareness stage.
Figure 2: Flow of harmful queries in WildJailbreak through Qwen3-32B, showing the relationship between the final response (Answer/Refuse) and risk awareness during reasoning (Yes/No).
Main Cause of Safety Failure
Based on the results of Figure 2, we categorize safety failures into two distinct patterns: (1) Harm Misidentification, where the model fails to recognize harmful intent, and (2) Mismatch between risk awareness and response, where the model explicitly identifies risks but produces unsafe outputs. While the former indicates a detection failure, the latter reveals a more critical flaw in the model’s reasoning process.
Further investigation of the second pattern reveals a pattern in which the model’s subsequent reasoning effectively revises or overrides its own safety assessment, leading to unsafe outputs even after risks have been identified. We refer to this phenomenon as Self-Jailbreak.
To determine whether Self-Jailbreak is an isolated anomaly or a widespread issue, we extended our investigation across multiple Large Reasoning Models (LRMs), as shown in Figure 3. Our results yield a significant finding:
Self-Jailbreak constitutes the predominant safety failure mode across LRMs. As illustrated in Figure 2, Self-Jailbreak consistently surpasses Harm Misidentification as the primary source of unsafe outputs across diverse model families and parameters. For instance, in DeepSeek-R1, 93.7% of safety failures are attributable to Self-Jailbreak. This indicates that Self-Jailbreak represents a systemic and persistent challenge, even for state-of-the-art LRMs.
Figure 3: Proportion of the safety failure cause across multiple LRMs conditioned on the unsafe responses of Wildjailbreak
Model Benign Reframing Warning Logical Fallacies
DS-R1 40.91 56.97 2.12
DS-Llama-70B 39.17 58.53 2.30
Qwen3-8B 36.94 57.91 5.15
Qwen3-14B 38.60 58.20 3.20
Qwen3-32B 36.86 58.96 4.18
Table 1: Distribution of self-jailbreak categories (%), where DS-R1 stands for DeepSeek-R1 while DS-Llama-70B represents DeepSeek-R1-Distill-Llama-70B.
2.2 A Taxonomy of Self-Jailbreak Behaviors
While the prevalence of self-jailbreak is evident, the underlying reasons that lead models to generate unsafe content remain poorly understood. A deeper analysis of how models override their own safety judgment is essential for designing effective safeguards. To this end, we conducted a qualitative analysis involving manual inspection of reasoning traces (Chain-of-Thought) and final responses, focusing on the discrepancy between the model’s internal risk analysis and its final output generation.
Taxonomy Definition
Based on recurring patterns identified in the manual inspection, we establish a taxonomy categorizing Self-Jailbreak into three distinct behaviors:
  • Benign Reframing: The model actively reinterprets the user’s malicious intent as benign (e.g., educational or theoretical), thereby justifying a helpful response.
  • Warning: The model assumes that appending a safety warning or disclaimer is sufficient to mitigate the harm, leading to a “warn-but-answer” failure mode.
  • Logical Fallacies: The model’s reasoning becomes entangled in complex or contradictory logical constraints within the prompt, causing it to bypass safety guardrails due to erroneous logical deductions (examples in Appendix E).
Figure 4: Overview of the Chain-of-Guardrail (CoG) framework. Phase 1: the original LRM produces an initial COT, which is decomposed into three atomic components—risk awareness $d(x)$, risk analysis $a(x)$, and response strategy p(x)p(x)—and classified for self-jailbreak risks. Phase 2: guided by the classification signal, CoG applies Safety Recomposition (SafR) or Safety Backtrack (SafB) to construct a safety-oriented COT (S-COT). Phase 3: The S-COT guides the model to generate the final safe response.
Quantitative Analysis
Utilizing this taxonomy, we annotated Self-Jailbreak instances across multiple LRMs to analyze their distribution. As shown in Table 1, a key finding emerges. Most notably, Warning is the primary Self-Jailbreak type, consistently accounting for over 55% of cases across all evaluated models. This suggests that the primary safety bottleneck is not a lack of risk awareness, but rather a flaw in the risk analysis. Models have learned to be overly compliant and often default to answering even when risks are identified, rather than refusing. The prevalent Warning and Benign Framing behavior indicates that during training, models may not have internalized clear signals for when to withhold responses entirely.
Implications for Methodology
Overall, these results indicate that Self-Jailbreak is both widespread and structured. Unsafe behavior often arises after correct risk recognition and recurs in a small number of type-specific reasoning patterns. This implies that coarse-grained safety enforcement or globally rewriting reasoning traces may discard valid reasoning steps while still missing the failure source, motivating mitigation that first identifies the Self-Jailbreak type and then targets the corresponding failure-inducing steps.
3 Chain-of-Guardrails: A Training Framework for Reasoning-Aware Safety
Motivated by the observation that safety failures in LRMs frequently arise during reasoning despite correct risk awareness, we propose Chain-of-Guardrails (CoG), a training framework that identifies and corrects unsafe reasoning steps while preserving the model’s inherent reasoning capability. Rather than imposing uniform safety constraints on the entire response, CoG performs targeted, reasoning-level interventions guided by diagnosed Self-Jailbreak behaviors.
3.1 Framework Overview
Given a query $x$ and an original LRM $\pi_{0}$, CoG aims to produce a safe response $y_{\text{safe}}$ while preserving the model’s original reasoning ability. To enable fine-grained analysis and intervention, we explicitly decompose the model’s reasoning trajectory into interpretable components.
3.2 Phase 1: Input & Analysis
In Phase 1, the original model $\pi_{0}$ generates an initial reasoning trajectory and response to the input query $x$. We decompose the resulting reasoning trajectory into three interpretable components:
$c=\pi_{0}(x)=[d(x),a(x),p(x)]$ (1)
where $d(x)$ denotes the model’s risk awareness, $a(x)$ its risk analysis, and $p(x)$ its response strategy.
Given the decomposed components, we run a Self-Jailbreak classifier to predict whether the reasoning trajectory contains Self-Jailbreak behaviors and, when applicable, output the corresponding Self-Jailbreak type.
3.3 Phase 2: Safety-Oriented Reasoning Transformation
In Phase 2, CoG transforms unsafe reasoning trajectories into safety-oriented ones by applying targeted interventions conditioned on the classification signal from Phase 1.
We introduce the following complementary transformation strategies:
@@ 3.3 Phase 2: Safety-Oriented Reasoning Transformation
Safety Backtrack (SafB) SafB keeps the original reasoning chain and adds a targeted self-check step. Guided by the Self-Jailbreak type, the original model $\pi_{0}$ takes the original $a(x)$ and $p(x)$ and generates a self-check segment that focuses on the failure-inducing parts of the reasoning. This self-check revisits the earlier reasoning decisions and provides corrective guidance before producing the final response. We then append the self-check segment to the end of the original chain, forming an augmented, safety-oriented reasoning trajectory.
3.4 Phase 3: Inference & Alignment
In Phase 3, the safety-oriented COT produced by SafR or SafB is used to generate a final safe response, which is then used as training data for alignment. We then fine-tune the original model $\pi_{0}$ on this constructed data using selective loss masking.
For SafR, we supervise the entire recomposed reasoning chain together with the final response. For SafB, we supervise only the appended self-check segment and the final response, while masking the original reasoning tokens. This selective supervision encourages learning safety-oriented corrections while mitigating unintended distributional shift away from $\pi_{0}$’s original reasoning trajectory.